OpenStudy (anonymous):

∫〖(5x^4 ln⁡〖x^5)dx〗 〗 evaluate in parts

OpenStudy (anonymous):

\[ ∫〖(5x^4 \ln⁡〖x^5)dx〗 〗\]

OpenStudy (anonymous):

\[ ∫▒〖(5x^4 \ln⁡x^5)dx〗 〗\]

OpenStudy (anonymous):

evaluation in parts is key here is an example

OpenStudy (anonymous):

\[ \begin{eqnarray*} \int 5x^4 \ln(x^5)\ dx &=& 5\int x^4 \ln(x^5)\ dx \\&=& \ln(x^5)x^5 - 5\int x^4\ dx \\ &=& \ln(x^5)x^5 - x^5 \end{eqnarray*} \]

OpenStudy (anonymous):

x/4 + 1/12x^3

OpenStudy (anonymous):

Wolfram agrees with me: http://www.wolframalpha.com/input/?i=integral+of+5x^4*ln%28x^5%29

OpenStudy (anonymous):

one sec i will post a very close example and you guys see if you can write it out in that format

OpenStudy (anonymous):

From Mathematica 8 Home Edition:\[\int\limits 5x^4\text{Log}\left[x^5\right]dx=5 \left(-\frac{x^5}{5}+\frac{1}{5} x^5 \text{Log}\left[x^5\right]\right) \]\[\text{Simplify}\left[5 \left(-\frac{x^5}{5}+\frac{1}{5} x^5 \text{Log}\left[x^5\right]\right)\right]=x^5 \left(-1+\text{Log}\left[x^5\right]\right)=x^5 \log \left(x^5\right)-x^5 \]